OpenStudy (anonymous):

If I wanted to create a formula to show an additional increase to something, would I use exponential or linear form? For example: A town produces enough food to feed 12 million people, and is increasing its food supply to feed an additional 160,000 people per year. How would I show that they've increased their current food supply to feed an additional 2 million people?

OpenStudy (anonymous):

add implies linear

OpenStudy (anonymous):

Thanks... So could I write my formula as: y= b+mx let b = 12000000+2,000,000 let m = 160000 let x = year y= 14,000,000 +160,000x ?